OpenStudy (anonymous):

11^7 / 11^6 simplify using the rules of exponets?

OpenStudy (aum):

\(\Large \frac{a^m}{a^n} = a^{m-n} \)

OpenStudy (asevilla5):

Step by step solution : Step 1 : 117 Simplify ——— 116 Dividing exponents : 1.1 117 divided by 116 = 11(7 - 6) = 111 = 11 Final result : 11

OpenStudy (asevilla5):

the simplify part is supposed to be 11^7/11^6 my bad

OpenStudy (aum):

\(\Large \frac{a^m}{a^n} = a^{m-n}\) \(\Large \frac{11^7}{11^6} = 11^{7-6} = 11^1 = 11\)

OpenStudy (anonymous):

Thank you